OpenStudy (anonymous):

What is the slope of the graph of 2x – 6y = 9? A. 1/3 B. -1/3 C. -3 D. 3

OpenStudy (mathmale):

Have you tried solving 2x – 6y = 9 for y? Can you think of any reason for doing that?

OpenStudy (mathmale):

Where have you seen y = mx + b before, and what does this represent?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

its A y=mx+c u can compare from here

OpenStudy (anonymous):

thanks @ikatouni

OpenStudy (isaiah.feynman):

@ikatouni You will have to get your equation into the slope-intercept form in order to see the slope. You do that by solving for y in that equation.